Suppose the risk-free interest rate is 2% and the market risk premium is 6%. A company, ABC, has a beta of 1.2. The Dividend per share of $1.1 was just paid to the investors. The dividend growth is 6% per year in the next two years and 5% per year for all years after that. The retention percentage rate is 50%. The stock is fairly priced.

(a) Calculate the present value of growth opportunities for ABC.

(b) Would this answer change if the stock was not fairly priced? (Explanation & Workings if Yes.)
Answer #1

Calculating Cost of Equity,

Using TVM Calculation,

Cost of Equity = 0.02 + 1.20(0.06)

Cost of Equity = 9.20%

Stock Price = 1.10(1.06)/1.092 + 1.10(1.06)2/(1.092)2 + 1.10(1.06)2(1.05)/(0.092 - 0.05)(1.092)2

Stock Price = $28.02

PVGO = Stock Price - Earning/Cost of equity

PVGO = 28.02 - 2.332/0.092

PVGO = $2.67